# Isotopy classification of Morse polynomials of degree three on ${\mathbb{R}}^{3}$

### v8

@@ 2.2. Up-down involution
Proof (see Fig. 2). Let $c_{i}$, $i\in\{1,\dots,r\}$, be a real critical value of $f$. Together with the vanishing cycle $\Delta_{i}(c_{i}-\varepsilon)\subset f^{-1}(c_{i}-\varepsilon)$, the cycle $\Delta^{+}_{i}(c_{i}+\varepsilon)\subset f^{-1}(c_{i}+\varepsilon)$ can be defined in the same way. It is easy to compute (see Lemma 1 in §V.3.1 of [28]) that the transport over the arc $\{c_{i}+\varepsilon e^{it},\},t\in[0,\pi],$ moves one of these two cycles to the other multiplied by $(-1)^{k}$, where $k$ is the positive inertia index of the $i$-th critical point, see (4). By the construction, the central symmetry in ${\mathbb{C}}^{3}$ geometrically moves the cycle in $f^{-1}(0)$, which is transported from the cycle $\Delta^{+}_{i}(c_{i}+\varepsilon)$ over a path in the upper half-plane, into a cycle in $\check{f}^{-1}(0)$, which is transported from the cycle $\tilde{\Delta}_{r+1-i}(-c_{i}-\varepsilon)\subset\check{f}^{-1}(-c_{i}-\varepsilon)$ over a path in the lower half-plane. Moreover, this move preserves the canonical orientations of these cycles (in fact, one of them is multiplied by $(-1)^{3}\times(-1)\equiv 1,$ where the first factor means the change of the orientation of ${\mathbb{R}}^{3}$ at the central symmetry, and the second factor results from a comparison of the directions of the gradients of our functions). The complex conjugation moves this cycle in $\check{f}^{-1}(0)$ to the cycle transported over a path in the upper half-plane from the cycle complex conjugate to $\tilde{\Delta}_{r+1-i}(-c_{i}-\varepsilon)$. The latter cycle is equal to the cycle $\tilde{\Delta}_{r+1-i}(-c_{i}-\varepsilon)$ itself multiplied by $(-1)^{3-k}$ (where 3k3-k is the positive inertia index of the critical point of $\check{f}$ with critical value $c_{i}$). This proves the statement of our proposition about cycles vanishing in real critical points. The statement about the non-real ones is obvious. (In fact, the exact value of the sign $\pm$ in it is also determined by the virtual Morse function of $f$, but its description is too long). $\Box$
@@ 2.3. D-graph invariant
The D-graph of a generic real Morse polynomial $f$ with only real critical points is [the isomorphism class of] the oriented graph with vertices of two types (“even” and “odd”), obtained from the Coxeter-Dynkin graph of $f$ by
@@
b)) an integer number in the range from 0 to 88 ((describing the last element of the virtual Morse function, i.e., the number of negative critical values)). $\Box\ \Box$
@@ 2.4. Independent perturbations of critical points
A neighborhood of a polynomial of degree 3 with a non-discriminantal principal homogeneous part in the space of all such polynomials can be considered as a piece of the parameter space of a versal deformation of the polynomial given by this principal part.
@@ 3. Fundamental lemma
Suppose that $f$ is a Morse polynomial whose four critical points with one of these two index collections lie in the plane $\{z=0\}$. Let
@@
The homogeneous principal part of the polynomial $\varphi$ is not discriminantal ((i.e., it vanishes on three different lines in ${\mathbb{C}}^{2})$.
Proof. If this principal part is discriminantal but $\varphi$ has only isolated critical points in ${\mathbb{C}}^{2}$, then by the Bezout’s theorem the number of these points is less than 4. If $\varphi$ has non-isolated critical points but $\varphi\not\equiv 0$ then all its real critical points form a line in ${\mathbb{R}}^{2}$. The critical points of the entire function $f$ lying in ${\mathbb{R}}^{2}$ are the intersection points of this line with the zero set of the quadratic polynomial $f_{2}$. This intersection set is either infinite (in which case the entire function $f$ has infinitely many critical points) or consists of at most two points. Finally, if $\varphi\equiv 0$ then $f$ has non-isolated critical points in ${\mathbb{C}}^{3}$. $\Box$
@@
If a polynomial $\varphi:{\mathbb{R}}^{2}\to{\mathbb{R}}$ of degree three with non-discriminantal principal homogeneous part has four real Morse critical points, then one of two things is true:
@@ 4.1. Polynomials with eight critical points
Figure 6. $D_{4}^{+}+D_{4}^{+}$ and $D_{5}+A_{3}$, and the last side of $D_{4}^{-}+D_{4}^{-}$; Card = 72
Figure 7. Other side of $D_{4}^{-}+D_{4}^{-}$, and also of $E_{6}+A_{2}$; Card = 432
@@
Proof. This follows easily from Proposition 11, cf. Example on p. 11. $\Box$
@@
has two critical points of the class $D_{4}^{-}$ (see Table 1 on page 1). In a neighborhood of one of them this polynomial is equivalent to
@@
Each of these two critical points can be further split into four Morse critical points by a small perturbation. Namely, the polynomial
@@
has one point of signature $(3,0)$, four points of signature $(2,1)$ and three points of signature $(1,2)$. So, we get Morse polynomials with passports $(0,4,4,0)$, $(1,3,3,1)$, $(0,3,4,1)$, and $(1,4,3,0)$.
@@
The D-graphs of these four polynomials with passports $(0,4,4,0)$, $(1,3,3,1)$, $(0,3,4,1)$, $(1,4,3,0)$ are respectively the D-graphs shown in Figs. 5, 6, 7, and the graph obtained from Fig. 7 by the up-down involution described in Proposition 23.
Proof. In all cases, the matrix of the intersection indices of four vanishing cycles arising from either of the critical points (16) or (17) is represented by the standard Coxeter–Dynkin graph of type $D_{4}$. Therefore, for each of the four polynomials (18)–(21), we can divide the set of vertices of the corresponding D-graph into two subsets of cardinality four in such a way that, by removing all edges connecting vertices from different subsets, we split the D-graph into two such graphs (with some orientation of the edges and coloring of the vertices depending on the choice of perturbations). The D-graphs of Figs. 8 and 9 and their images under the up-down involution do not have such splittings.
For each of the D-graphs of Figs. 5, 6, 7 and the up-down reflection of Fig. 7, the Morse indices of the corresponding critical points are given in Proposition 31. Their passports fix the correspondence between these D-graphs and four polynomials (18)–(21). $\Box$
@@
The corresponding D-graph should contain two subgraphs with complementary sets of 2 and 6 vertices, which are isomorphic to classical Coxeter–Dynkin graphs of types $A_{2}$ and $E_{6}$, and all edges of this D-graph connecting these subgraphs are directed from the $E_{6}$ to the $A_{2}$ graph. Of all the D-graphs mentioned in Proposition 29 (including the up-down versions of the graphs explicitly drawn there), only the one from Fig. 8 satisfies this property. So, the polynomial (23) realizes exactly this D-graph, and minus this polynomial realizes its up-down version.
@@
has two real critical points of class $A_{4}$, see § 7.3 of the paper [32]. Let us perturb these critical points so that the resulting Morse function has only real critical points. The D-graph of this function can be split into two subgraphs of type $A_{4}$ in such a way that all its edges connecting these subgraphs are oriented from one subgraph to the other. Only one of D-graphs 59 admits such splitting. The polynomial (25) is of class $\Xi_{2}$. Indeed, each its $A_{4}$ singularity can be perturbed so that the resulting function has no real critical points. By §1.2, this is impossible for polynomials of type $\Xi_{2}$.
@@ 4.1.4. Proof of Theorem 12 for polynomials of type $\Xi_{1}$
1. The set $\{E_{6}+A_{2}\}$ of all third degree polynomials ${\mathbb{R}}^{3}\to{\mathbb{R}}$ with non-discriminantal principal homogeneous parts and two real critical points of classes $E_{6}$ and $A_{2}$ consists of exactly four smooth connected strata. Every polynomial with these properties can be reduced by an orientation preserving affine transformation of ${\mathbb{R}}^{3}$ and adding a constant to one of the normal forms
@@
Thus, for the parameters of the intersection surfaces we can in all cases take the coefficient $\varkappa\in(-\infty,0)$ or $(0,+\infty)$ and the constant that can be added to the function ((27)) or ((28).).
@@
Let $f_{-}$ and $\tilde{f}_{-}$ be near endpoints of the paths $I$ and $\tilde{I}$ that lie in this neighborhood of $f!$ and have equal sets of critical values. By the construction, the virtual Morse functions associated with them are the same, in particular their two Coxeter–Dynkin graphs are the same and contain equal subgraphs with complementary sets of 6 and 2 vertices corresponding to the critical points of $f_{-}$ and $\tilde{f}_{-}$ obtained by decomposing two critical points of $f!$. Consider the projections of the points $f_{-}$ and $\tilde{f}_{-}$ to the parameter space of the standard miniversal deformation
@@
In the first case (when $f_{-}=\tilde{f}_{-}$) we continue this equality along our paths and have $f=\tilde{f}$. In the second case, the identity $\tilde{f}_{\tau}(x,y,z)\equiv f_{\tau}(z,y,x)$ holds by the continuity over our paths for all their points, including the starting points $f$ and $\tilde{f}$, because the Lyashko–Looijenga covering commutes with the involution $x\leftrightarrow z$. $\Box$
@@
2. The D-graph of Fig. 9 has no non-trivial symmetries preserving the colors of the vertices, so all four critical points with Morse index 1 of any Morse polynomial with this D-graph can be ordered in an invariant way depending continuously of these polynomials. By Lemma 24 these four ordered points never lie in the same affine plane, and thus define an orientation of ${\mathbb{R}}^{3}$. The polynomials of the same isotopy class define equal orientations, but any polynomial and its reflection define opposite orientations. $\Box$
@@ 4.2. Polynomials with six real critical points
The set of all 1512 virtual Morse functions of type $\Xi_{1}$ with exactly six real critical points splits into five virtual components: one of cardinality 28 and passport $(1,2,2,1)$, two of cardinality 112 and passports $(1,3,2,0)$ and $(0,2,3,1)$, and two of cardinality 630 and passport $(0,3,3,0)$.
@@
is invariant under the reflection in the plane $\{y=0\}$, so its isotopy class is also achiral.
@@ 4.3. Polynomials with four critical points
The set of all 515 virtual Morse functions of type $\Xi_{1}$ with exactly four real critical points splits into four virtual components: one of cardinality 30, two of cardinality 60, and one of cardinality 365.
Proof. As before, our program finds three virtual components with Card equal to 30, 60 and 365. A component with Card=60 contains the virtual Morse function (35) (left).
@@ 4.4. Polynomials with two critical points
The set of all 390 virtual Morse functions of type $\Xi_{1}$ with exactly two real critical points splits into three virtual components: two of cardinality 102 and one of cardinality 186.
@@
These components are represented by the polynomial
@@
for passport $(1,1,0,0)$, minus this polynomial for passport $(0,0,1,1)$, and the polynomial
@@
with sufficiently small $\varepsilon>0$ for passport $(0,1,1,0)$. In the latter case, the additional term $\varepsilon x(1+2z)$ moves all critical point of the polynomial (22) that are close to the origin into the non-real domain, and splits its critical point of class $A_{2}$ into two real points with Morse indices 1 and 2.
@@ 5.1. Polynomials with eight critical points
of type $\Xi_{2}$. It has a critical point of class $E_{6}$ at the origin, and additionally two Morse critical points with Morse indices 1 and 2. The critical values at the last two points are equal to each other and to $\gamma^{3}$. For $\gamma>0$ this polynomial has the form $u^{3}-v^{4}+w^{2}$ in some local coordinates at the origin. This its critical point can be split into three Morse points with Morse index 1 and three points with index 2 (see § 4.1.2). The intersection matrix of the vanishing cycles of the corresponding Morse polynomial is computed in § 7 of [29] and indeed defines the D-graph of Fig. 13.
@@
The substitution $X=\tilde{x}-\frac{4}{3}\tilde{z}$ gives the form
@@
On the other hand, the Milnor number of $f$ at the origin is at most 5, since the sum of the Milnor numbers of all critical points in ${\mathbb{C}}^{3}$ of any polynomial of type $\Xi_{2}$ is equal to eight. The rank of this critical point is obviously equal to two, so this is a point of class $A_{5}$. The critical points obtained by any its morsification can only have Morse indices 1 and 2 because the quadratic part of $f$ has signature $(1,1)$. The number of all critical points with even Morse indices of any morsification of $f$ is equal to the number of critical points with odd indices, so any decomposition of this $A_{5}$ singularity into five Morse critical points has three points with Morse index 1 and two points with index 2.
@@
1. The set of all third degree polynomials ${\mathbb{R}}^{3}\to{\mathbb{R}}$ with non-discriminantal principal homogeneous parts and three critical points of classes $E_{6}$, $A_{1}$, and $A_{1}$ with coinciding critical values at two $A_{1}$ points consists of exactly two smooth connected strata of codimension 6 in the space of all third degree polynomials. All these polynomials can be reduced by orientation preserving affine transformations of ${\mathbb{R}}^{3}$ and by adding constants to the normal form ((43)) with $\gamma\neq 0$ ((and the strata are characterized by the sign of $\gamma)$. In particular, all such polynomials are of type $\Xi_{2}$. The $j$-invariant of the principal homogeneous parts of all of them is equal to 1728.
2. The intersection of each of these two strata with the space $\Theta\cap{\mathbb{R}}^{8}$ of real polynomials ((7)) with $A\neq 1$ is the union of three smooth surfaces diffeomorphic to ${\mathbb{R}}^{2}$, one of which consists of all polynomials
@@
where the sign of the coefficient $\varkappa\neq 0$ matches the sign of the coefficient $\gamma$ in ((43)) that characterizes the stratum, $X\equiv x-\varkappa$, $Y\equiv y-(4+2\sqrt{3})\varkappa$ and $Z\equiv z-\varkappa$, and two other surfaces are obtained from this one by permutations of the coordinates.
@@
The $j$-invariant of the principal part of (43) is equal to 1728. Conversely, a polynomial in the Weierstrass normal form (2) has $j=1728$ if and only if $b=0,a\neq 0$. If $a>0$ then it is of type $\Xi_{1}$, and if $a<0$ then it can be reduced to $x^{3}-xz^{2}-y^{2}z$ by dilating the coordinates.
@@
3. For any $t>0$, the action
(46) $f_{A,\lambda}(x,y,z)\mapsto t^{3}f_{A,\lambda}(t^{-1}x,t^{-1}y,t^{-1}z)\hskip 85.35826pt$
preserves the principal homogeneous part of the polynomial, multiplies the coefficient $\varkappa$ by $t$ and multiplies all critical values by $t^{3}$. $\Box$
@@ 3. For any $t>0$, the action
Let $f$ and $\tilde{f}$ be two generic polynomials of type $\Xi_{2}$ with equal virtual Morse functions. Arguing as in § 4.1.4, we see that there exist generic polynomials $g$ and $\tilde{g}$ of the form (7) such that a) $g$ is isotopic to $f$ and $\tilde{g}$ is isotopic to $\tilde{f}$ in the class of generic polynomials, b) $g$ and $\tilde{g}$ have equal sets of critical values and standard paths in ${\mathbb{C}}^{1}$ defining the vanishing cycles, and c) there are two paths connecting $g$ and $\tilde{g}$ to $f!$ in the space of Morse polynomials of type $\Xi_{2}$ and form (7) such that the images of these two paths under the Lyashko–Looijenga map coincide and their near endpoints are polynomials with the D-graph of Fig. 13. It follows from the structure of a neighborhood of the point $f!$ in $\Theta$ that these near endpoints either coincide with each other (and then, by continuity along our paths, so do the starting points $g$ and $\tilde{g}$), or they are reduced to each other by the unique involution of this neighborhood preserving the Lyashko–Looijenga map: this involution is induced by the reflection $(x,y,z)\leftrightarrow(z,y,x)$ of ${\mathbb{R}}^{3}$. In the latter case, our paths are mapped to each other by this involution, in particular their starting polynomials $g$ and $\tilde{g}$ are also mapped to each other by the reflection of arguments in the plane $\{x=z\}$. $\Box$

### v9

@@ 2.3. Up-down involution
Proof (see Fig. 2). Let $c_{i}$, $i\in\{1,\dots,r\}$, be a real critical value of $f$. Together with the vanishing cycle $\Delta_{i}(c_{i}-\varepsilon)\subset f^{-1}(c_{i}-\varepsilon)$, the cycle $\Delta^{+}_{i}(c_{i}+\varepsilon)\subset f^{-1}(c_{i}+\varepsilon)$ can be defined in the same way. It is easy to compute (see Lemma 1 in §V.3.1 of [33]) that the transport over the arc $\{c_{i}+\varepsilon e^{it},\},t\in[0,\pi],$ moves one of these two cycles to the other multiplied by $(-1)^{k}$, where $k$ is the positive inertia index of the second differential of the $i$-th critical point, see (4). By the construction, the central symmetry in ${\mathbb{C}}^{3}$ geometrically moves the cycle in $f^{-1}(0)$, which is transported from the cycle $\Delta^{+}_{i}(c_{i}+\varepsilon)$ over a path in the upper half-plane, into a cycle in $\check{f}^{-1}(0)$, which is transported from the cycle $\tilde{\Delta}_{r+1-i}(-c_{i}-\varepsilon)\subset\check{f}^{-1}(-c_{i}-\varepsilon)$ over a path in the lower half-plane. Moreover, this move preserves the canonical orientations of these cycles (in fact, one of them is multiplied by $(-1)^{3}\times(-1)\equiv 1,$ where the first factor means the change of the orientation of ${\mathbb{R}}^{3}$ at the central symmetry, and the second factor results from a comparison of the directions of the gradients of our functions). The complex conjugation moves this cycle in $\check{f}^{-1}(0)$ to the cycle transported over a path in the upper half-plane from the cycle complex conjugate to $\tilde{\Delta}_{r+1-i}(-c_{i}-\varepsilon)$. The latter cycle is equal to the cycle $\tilde{\Delta}_{r+1-i}(-c_{i}-\varepsilon)$ itself multiplied by $(-1)^{3-k}$ (where 3k3-k is the positive inertia index of the critical point of $\check{f}$ with critical value $c_{i}$). This proves the statement of our proposition about cycles vanishing in real critical points. The statement about the non-real ones is obvious. (In fact, the exact value of the sign $\pm$ in it is also determined by the virtual Morse function of $f$, but its description is too long). $\Box$
@@ 2.4. D-graph invariant
The D-graph of a generic real Morse polynomial $f$ with only real critical points is <<the isomorphism class of>> the oriented graph with vertices of two types (“even” and “odd”), obtained from the Coxeter-Dynkin graph of $f$ by
@@
b)) an integer number in the range from 0 to 88 ((describing the last element of the virtual Morse function, i.e., the number of negative critical values)). $\Box$
Proposition 23 follows immediately from Lemmas 24 and 25. $\Box$
@@ 2.5. Independent perturbations of critical points
A neighborhood of a polynomial of degree three with a non-discriminant principal homogeneous part in the space of all such polynomials can be considered as a piece of the parameter space of a versal deformation of the polynomial given by this principal part.
@@ 3. Fundamental lemma
Suppose that $f$ is a polynomial whose four Morse critical points with one of these two index collections lie in the plane $\{z=0\}$. Let
@@
The homogeneous principal part of the polynomial $\varphi$ is not discriminant ((i.e., it vanishes on three different lines in ${\mathbb{C}}^{2})$.
Proof. If this principal part is discriminant but $\varphi$ has only isolated critical points in ${\mathbb{C}}^{2}$, then by the Bezout’s theorem the number of these points is less than 4. If $\varphi$ has non-isolated critical points but $\varphi\not\equiv 0$ then all its real critical points form a line in ${\mathbb{R}}^{2}$. The critical points of the entire function $f$ lying in ${\mathbb{R}}^{2}$ are the intersection points of this line with the zero set of the quadratic polynomial $f_{2}$. This intersection set is either infinite (in which case the entire function $f$ has infinitely many critical points) or consists of at most two points. Finally, if $\varphi\equiv 0$ then the polynomial $f$ is reducible and has non-isolated critical points in ${\mathbb{C}}^{3}$. $\Box$
@@
If a polynomial $\varphi:{\mathbb{R}}^{2}\to{\mathbb{R}}$ of degree three with non-discriminant principal homogeneous part has four real Morse critical points, then one of two things is true:
@@ 4.1. Polynomials with eight critical points
Figure 6. $D_{4}^{+}+D_{4}^{+}$, $D_{5}+A_{3}$, and $D_{4}^{-}+D_{4}^{-}$; Card = 72
Figure 7. $D_{4}^{-}+D_{4}^{-}$, $E_{6}+A_{2}$, and $D_{5}+A_{3}$; Card = 432
@@
Proof. This follows easily from Proposition 14, cf. Example on p. 14. $\Box$
@@
has two critical points of the class $D_{4}^{-}$ (see Table 2 on page 2). In a neighborhood of one of them this polynomial is equivalent to
@@
Each of these two critical points can be further split into four Morse critical points by a small perturbation of the polynomial (15). Namely, the polynomial
@@
has one point of signature $(3,0)$, four points of signature $(2,1)$ and three points of signature $(1,2)$. So, the passports of the Morse polynomials (18), (19), (20), and (21) are $(0,4,4,0)$, $(1,3,3,1)$, $(0,3,4,1)$, and $(1,4,3,0),$ respectively.
@@
The D-graphs of these four polynomials are respectively the D-graphs shown in Figs. 5, 6, 7, and the graph obtained from Fig. 7 by the up-down involution described in Proposition 26.
Proof. In all cases, the matrix of the intersection indices of four vanishing cycles arising from either of the critical points (16) or (17) is represented by the standard Coxeter–Dynkin graph of type $D_{4}$. Therefore, for each of the four polynomials (18)–(21), we can divide the set of vertices of the corresponding D-graph into two subsets of cardinality four in such a way that, by removing all edges connecting vertices from different subsets, we split the D-graph into two such graphs (with some orientation of the edges and coloring of the vertices depending on the choice of perturbations). The D-graphs of Figs. 8 and 9 and their images under the up-down involution do not have such splittings. Thus, according to Proposition 32, the D-graphs of the polynomials (18)–(21) are among the remaining four D-graphs described in this proposition.
For each of the D-graphs of Figs. 5, 6, 7 and the up-down reflection of Fig. 7, the Morse indices of the critical points of all polynomials having these D-graphs are determined by Proposition 34. In particular, the passports of these polynomials are equal to $(0,4,4,0)$, $(1,3,3,1)$, $(0,3,4,1)$, and $(1,4,3,0)$ respectively. This fixes the only possible correspondence between the polynomials (18)–(21) and these four D-graphs. $\Box$
@@
The corresponding D-graph should contain two subgraphs with complementary sets of 2 and 6 vertices, which are isomorphic to classical Coxeter–Dynkin graphs of types $A_{2}$ and $E_{6}$, and all edges of this D-graph connecting these subgraphs are directed from the $E_{6}$ graph to the $A_{2}$ graph. Of all the D-graphs mentioned in Proposition 32 (including the up-down versions of the graphs explicitly drawn there), only the one from Fig. 8 satisfies this property. So, the polynomial (23) realizes exactly this D-graph, and minus this polynomial realizes its up-down version.
@@
has two real critical points $(0,0,0)$ and $\left(\frac{1}{64},\frac{-1}{128},\frac{-7}{192}\right)$ of class $A_{4}$, see § 7.3 of the paper [38]. Let us perturb these critical points so that the resulting Morse function has only real critical points. The D-graph of this function can be split into two subgraphs of type $A_{4}$ in such a way that all its edges connecting these subgraphs are oriented from one subgraph to the other. Only one of the D-graphs 59 admits such splitting. The polynomial (25) is of class $\Xi_{1}$. Indeed, each its $A_{4}$ singularity can be perturbed so that the resulting function has no real critical points. According to §1.4, this is impossible for polynomials of type $\Xi_{2}$.
@@ 4.1.4. Proof of Theorem 15 for polynomials of type $\Xi_{1}$
1. The set $\{E_{6}+A_{2}\}$ of all third degree polynomials ${\mathbb{R}}^{3}\to{\mathbb{R}}$ with non-discriminant principal homogeneous parts and two real critical points of classes $E_{6}$ and $A_{2}$ consists of exactly four smooth connected strata. Every polynomial with these properties can be reduced by an orientation preserving affine transformation of ${\mathbb{R}}^{3}$ and adding a constant to one of the normal forms
@@
Thus, for the parameters of the intersection surfaces we can in all cases take the coefficient $\varkappa\in(-\infty,0)$ or $(0,+\infty)$ and the constant that can be added to the polynomial ((27)) or ((28).).
@@
Let $f_{-}$ and $\tilde{f}_{-}$ be near endpoints of the paths $I$ and $\tilde{I}$ that lie in this neighborhood of $f!$ and have equal sets of critical values. By the construction, the virtual Morse functions associated with them are the same. In particular, their two Coxeter–Dynkin graphs are the same and contain equal subgraphs with complementary sets of 6 and 2 vertices corresponding to the critical points of $f_{-}$ and $\tilde{f}_{-}$ obtained by decomposing two critical points of $f!$. Consider the projections of the points $f_{-}$ and $\tilde{f}_{-}$ to the parameter space of the standard miniversal deformation
@@
In the first case (when $f_{-}=\tilde{f}_{-}$) we continue this equality along our paths and have $f=\tilde{f}$. In the second case, the identity $\tilde{f}_{\tau}(x,y,z)\equiv f_{\tau}(z,y,x)$ holds by the continuity over our paths for all their points, including the starting points $f$ and $\tilde{f}$, because the Lyashko–Looijenga covering commutes with the involution induced by the reflection $x\leftrightarrow z$. $\Box$
@@
2. The D-graph of Fig. 9 has no non-trivial symmetries preserving the colors of the vertices, so all four critical points with Morse index 1 of any Morse polynomial with this D-graph can be ordered in an invariant way depending continuously of these polynomials. By Lemma 27 these four ordered points never lie in the same affine plane, and thus define an orientation of ${\mathbb{R}}^{3}$. The polynomials of the same isotopy class define equal orientations, but any polynomial and its mirror image define opposite orientations. $\Box$
@@ 4.2. Polynomials with six real critical points
The set of all 1512 virtual Morse functions of type $\Xi_{1}$ with exactly six real critical points splits into five virtual components: one with cardinality 28 and passport $(1,2,2,1)$, two with cardinality 112 and passports $(1,3,2,0)$ and $(0,2,3,1)$, and two with cardinality 630 and passport $(0,3,3,0)$. Each of these virtual components is achiral and so is associated with a single isotopy class of Morse polynomials.
@@
is invariant under the reflection in the plane $\{y=0\}$, so its isotopy class is also achiral. $\Box$
@@ 4.3. Polynomials with four critical points
The set of all 515 virtual Morse functions of type $\Xi_{1}$ with exactly four real critical points splits into four virtual components: one with cardinality 30, two with cardinality 60, and one with cardinality 365. Each of these virtual components is achiral and associated with a single isotopy class of Morse polynomials.
Proof. As before, our program finds three virtual components with Card equal to 30, 60 and 365. A virtual component with Card=60 contains the virtual Morse function (35) (left).
@@ 4.4. Polynomials with two critical points
The set of all 390 virtual Morse functions of type $\Xi_{1}$ with exactly two real critical points splits into three virtual components: two with cardinality 102 and one with cardinality 186. Each of these virtual components is associated with a single isotopy class of Morse polynomials.
@@
The passport of the polynomial
@@
is equal to $(1,1,0,0)$, the passport of minus this polynomial is equal to $(0,0,1,1)$, and the passport of the polynomial
@@
with sufficiently small $\varepsilon>0$ is equal to $(0,1,1,0)$. In the last case, adding the term $\varepsilon x(1+2z)$ to polynomial (22) moves all of its critical points that are close to the origin into the non-real domain. It also splits the critical point of class $A_{2}$ into two real points with Morse indices 1 and 2.
@@ 5.1. Polynomials with eight critical points
of type $\Xi_{2}$. It has a critical point of class $E_{6}$ at the origin, and additionally two Morse critical points with Morse indices 1 and 2. The critical values at the last two points are equal to each other and to $\gamma^{3}$. For $\gamma>0$ this polynomial has the form $u^{3}-v^{4}+w^{2}$ in some local coordinates at the origin. This its critical point can be split into three Morse points with Morse index 1 and three points with Morse index 2 (see § 4.1.2). The intersection matrix of the vanishing cycles of the corresponding Morse polynomial is computed in § 7 of [34] and indeed defines the D-graph of Fig. 13.
@@
The substitution $X=\tilde{x}-\frac{4}{3}\tilde{z}$ reduces it to the form
@@
On the other hand, the Milnor number of $f$ at the origin is at most 5, since the sum of the Milnor numbers of all critical points in ${\mathbb{C}}^{3}$ of any polynomial of type $\Xi_{2}$ is equal to eight. The signature of the second differential of this critical point is obviously equal to $(1,1)$. Therefore, this is a point of class $A_{5}$, and the critical points obtained by any its morsification can only have Morse indices 1 and 2. The number of all critical points with even Morse indices of any morsification of $f$ is equal to the number of critical points with odd indices, so any decomposition of this $A_{5}$ singularity into five Morse critical points has three points with Morse index 1 and two points with index 2.
@@
1. The set of all third degree polynomials ${\mathbb{R}}^{3}\to{\mathbb{R}}$ with non-discriminant principal homogeneous parts and three critical points of classes $E_{6}$, $A_{1}$, and $A_{1}$ with coinciding critical values at two $A_{1}$ points consists of exactly two smooth connected strata of codimension 6 in the space of all third degree polynomials. All these polynomials can be reduced to the normal form ((43)) with $\gamma\neq 0$ by orientation preserving affine transformations of ${\mathbb{R}}^{3}$ and by adding constants; the strata are characterized by the sign of $\gamma$. In particular, all such polynomials are of type $\Xi_{2}$. The $j$-invariant of the principal homogeneous parts of all these polynomials is equal to 1728.
2. The intersection of each of these two strata with the space $\Theta\cap{\mathbb{R}}^{8}$ of real polynomials ((7)) with $A\neq 1$ is the union of three smooth surfaces diffeomorphic to ${\mathbb{R}}^{2}$. One of these surfaces consists of all polynomials
@@
where the sign of the coefficient $\varkappa\neq 0$ matches the sign of the coefficient $\gamma$ in ((43)) that characterizes the stratum, $X\equiv x-\varkappa$, $Y\equiv y-(4+2\sqrt{3})\varkappa$ and $Z\equiv z-\varkappa$. Two other surfaces are obtained from this one by permutations of the coordinates.
@@
The $j$-invariant of the principal part of (43) is equal to 1728. Conversely, a polynomial in the Weierstrass normal form (2) has $j=1728$ if and only if $b=0,a\neq 0$. If $a>0$ then it is of type $\Xi_{1}$, and if $a<0$ then it can be reduced to the form $x^{3}-xz^{2}-y^{2}z$ by dilating the coordinates.
@@
3. The multiplicative group $\mathbb{R}$$_{+}$ of positive numbers acts on the space $\Xi_{2}$: for any $t>0$, the corresponding transformation is given by the formula
(46) $f_{A,\lambda}(x,y,z)\mapsto t^{3}f_{A,\lambda}(t^{-1}x,t^{-1}y,t^{-1}z).\hskip 85.35826pt$
This transformation preserves the principal homogeneous part of the polynomial, multiplies the coefficient $\varkappa$ by $t$ and multiplies all critical values by $t^{3}$. $\Box$
@@
Let $f$ and $\tilde{f}$ be two generic polynomials of type $\Xi_{2}$ with equal virtual Morse functions. Arguing as in § 4.1.4, we see that there exist generic polynomials $g$ and $\tilde{g}$ of the form (7) such that a) $g$ is isotopic to $f$ and $\tilde{g}$ is isotopic to $\tilde{f}$ in the class of generic polynomials, b) $g$ and $\tilde{g}$ have equal sets of critical values and standard paths in ${\mathbb{C}}^{1}$ defining the vanishing cycles, and c) there are two paths connecting $g$ and $\tilde{g}$ to $f!$ in the space of type $\Xi_{2}$ Morse polynomials of the form (7) such that the images of these two paths under the Lyashko–Looijenga map coincide and their near endpoints are polynomials with the D-graph of Fig. 13. It follows from the structure of a neighborhood of the point $f!$ in $\Theta$ that these near endpoints either coincide with each other (and then, by continuity along our paths, so do the starting points $g$ and $\tilde{g}$), or they are reduced to each other by the unique involution of this neighborhood preserving the Lyashko–Looijenga map: this involution is induced by the reflection $(x,y,z)\leftrightarrow(z,y,x)$ of ${\mathbb{R}}^{3}$. In the latter case, our paths are mapped to each other by this involution, in particular their starting polynomials $g$ and $\tilde{g}$ are also mapped to each other by the reflection of arguments in the plane $\{x=z\}$. $\Box$